in acertain country the true probablilty of a baby being a girl is 0.466. among the next nine randomly selected births in the country, what is the probability that at least one of them is a boy


1
Expert's answer
2021-09-23T16:42:31-0400

The probability of s baby being a boy is 1-0.466=0.534.

"P(X\ge1)=1-P(0)=1-C_9^00.534^00.466^9=1-0.0010=0.9990."